Find an expression in terms of (x) for 150% of 4y+6 if (x) is 40% of 2y+3.
I am so lost on this one.

Answer by math_il(6)

given that x = 40% of (2y+3)
expanding this
x = 40/100 * ( 2y+3)
40/100 can be simplified to 2/5 ( see note1)
x = 2/5 * ( 2y+3)
multiply by 5 on both sides
5x = 2*(2y+3)
multiply the terms on the right
5x = 4y + 6
So we get 4y +6 = 5x. ( A)
We want 150% of 4y+6
or 150% of 5x ( using A)
150% of 5x = 150/100 * 5x = 3/2( 5x) = 15/2 x. This is the answer.
note1: 40/100 = 4/10 = 2/5 ( divide by 2)